The surface area of a sphere inscribed in a regular tetrahedron is 144 cm2
22. What is the altitude of the tetrahedron?.

a) 20 cm
b) 24 cm
c) 28 cm
d) 48 cm
1
Expert's answer
2020-12-17T08:30:56-0500


Let "a=" base of a regular tetrahedron, "l=" slant height of a a regular tetrahedron, "r=" radius of a sphere inscribed in a regular tetrahedron.

The surface area of a sphere


"A=4\pi r^2"

Given "A=144cm^2"

"4\pi r^2=144""r^2=\dfrac{36}{\pi}"


"r=\dfrac{6\sqrt{\pi}}{\pi}\approx3.4(cm)""SD=l=a\dfrac{\sqrt{3}}{2}, ED=\dfrac{a}{2\sqrt{3}}"

From the right triangle "SED"


"SD^2=SE^2+ED^2"

"l^2=h^2+(\dfrac{a}{2\sqrt{3}})^2"

"\dfrac{3}{4}a^2=h^2+\dfrac{1}{12}a^2"

"h=\dfrac{\sqrt{6}}{3}a"

"\angle ESD=30\degree, \angle EDS=60\degree"

Right triangle "SOF"


"SO=2OF=2r"

"h=SE=SO+OE=2r+r=3r"

"h=3r=\dfrac{18\sqrt{\pi}}{\pi}(cm)\approx10.155(cm)"

The altitude of the tetrahedron is "h=\dfrac{18\sqrt{\pi}}{\pi}cm\approx10.155cm"

No right answer.